# Secure Digital Semantic Communications: Fundamentals, Challenges, and Opportunities

### v2

@@ Abstract
Semantic communication (SemCom) has emerged as a promising paradigm for future wireless networks by prioritizing task-relevant meaning over raw data delivery, thereby reducing communication overhead and improving efficiency. However, shifting from bit-accurate transmission to task-oriented delivery introduces new security and privacy risks. These include semantic leakage, semantic manipulation, knowledge base vulnerabilities, model-related attacks, and threats to authenticity and availability. Most existing secure SemCom studies focus on analog SemCom, where semantic features are mapped to continuous channel inputs. In contrast, digital SemCom transmits semantic information through discrete bits or symbols within practical transceiver pipelines, offering stronger compatibility with real-world systems while exposing a distinct and underexplored attack surface. In particular, digital SemCom typically represents semantic information over a finite alphabet through explicit digital modulation, following two main routes: probabilistic modulation and deterministic modulation. These discrete mechanisms and practical transmission procedures introduce additional vulnerabilities affecting bit- or symbol-level semantic information, the modulation stage, and packet-based delivery and protocol operations. Motivated by these challenges and the lack of a systematic analysis of secure digital SemCom, this paper reviews SemCom fundamentals, clarifies the architectural differences between analog and digital SemCom and their security implications, organizes the threat landscape for digital SemCom, and discusses potential defenses. Finally, we outline open research directions toward secure and deployable digital SemCom systems.
@@ I Introduction
However, the characteristics that make SemCom efficient and task-oriented also raise new concerns for security and privacy. Since SemCom often transmits task-oriented representations, the transmitted content may inadvertently reveal sensitive information. In particular, an eavesdropper may not need to reconstruct the source data to obtain useful information, because semantic information can already enable inference of private attributes. Beyond passive leakage, SemCom is also vulnerable to semantic manipulation, where an attacker perturbs the source or the semantic information to distort the recovered meaning and mislead downstream tasks, potentially causing misinformation, unsafe decisions, or service disruption. In addition, SemCom commonly relies on data-driven models and knowledge bases, expanding the attack surface beyond the physical channel. Adversaries may compromise the learning pipeline through data poisoning or backdoor insertion, exploit model-related vulnerabilities such as model extraction and inversion, or manipulate the knowledge base to bias semantic interpretation and induce systematic errors. Addressing these security and privacy challenges is therefore essential for practical deployment.
@@
In practice, analog and digital SemCom differ in how semantic information is transmitted over the wireless interface, leading to distinct attack surfaces and design constraints. Analog SemCom typically maps semantic features directly to continuous channel inputs and relies on joint source-channel behavior. Digital SemCom, in contrast, transmits semantic information through discrete bits or symbols within practical digital transceiver pipelines. In such designs, semantic information is mapped into a finite alphabet via explicit digital modulation, following two main routes: probabilistic modulation, which learns a transition distribution and generates discrete symbols using neural network models, and deterministic modulation, which integrates explicit quantization. Meanwhile, this discrete interface also facilitates the integration of cryptographic primitives. This pipeline introduces additional vulnerabilities, including bit- or symbol-level semantic leakage and manipulation. Moreover, digital SemCom also faces threats such as vulnerabilities within probabilistic or deterministic modulation components and attacks arising from packet-based delivery and protocol operations. These characteristics, together with the underexplored security risks in digital SemCom, motivate our dedicated focus on secure digital SemCom to provide a structured understanding of its threat landscape and corresponding defenses.
@@ II Background Knowledge of Semantic Communication
Figure 2: The detailed architecture of a digital SemCom system with explicit probabilistic or deterministic modulation and packet-based delivery. The JSC encoder/decoder denotes the joint source and channel coding/decoding module.
@@
The transmitter then modulates the semantic information into a transmitted signal, which propagates through the wireless channel and experiences fading, noise, and interference. At the receiver, the received signal is demodulated, and channel decoding mitigates channel distortions to recover reliable semantic information. The semantic decoder then reconstructs the intended meaning or directly outputs the task result for the information destination, which is defined at the receiver and reflects the target task. Throughout this process, a knowledge base can be maintained to store domain knowledge and contextual information, helping the transmitter and receiver align semantic interpretation and improve semantic consistency under limited communication resources.
@@
In probabilistic modulation [11], the modulator is modeled as a stochastic mapping implemented by a neural encoder that outputs a conditional distribution over a finite bit or symbol alphabet. The transmitted sequence is generated by sampling from this learned distribution, while the demodulator is realized by a probabilistic inference network that estimates the posterior distribution of the transmitted bits or symbols given the received signal. A key challenge is that end-to-end training must handle discrete random variables, making posterior inference and gradient estimation nontrivial. As a result, probabilistic digital SemCom often adopts techniques such as variational inference, score-function estimators, reparameterization-based approximations, and continuous relaxations such as Gumbel-Softmax, while accounting for potential mismatch between relaxed training and discrete inference at deployment.
@@ III-A Common Challenges in Analog and Digital Semantic Communication
Semantic leakage and inference attacks are among the most fundamental privacy challenges in SemCom. Because the transmitted message is task-oriented semantic information, intercepted semantic information can already reveal sensitive attributes, intents, or environmental states, especially when combined with side information or public background knowledge. SemCom is also vulnerable to semantic manipulation, where an attacker introduces carefully designed perturbations to distort the recovered meaning and mislead downstream tasks. Such attacks may not be reflected by conventional link-level metrics while still causing severe semantic errors. They also raise concerns about semantic integrity and source authenticity, since the receiver must determine whether the recovered meaning is reliable and originates from a legitimate transmitter rather than from fabricated or manipulated semantic information.
Another set of shared challenges arises from the supporting context and learning components. Many SemCom designs rely on shared context or knowledge base to align interpretation across endpoints, which introduces risks such as knowledge poisoning, manipulation, and inconsistencies between transmitter and receiver knowledge states that can systematically bias semantic interpretation. Moreover, because semantic extraction and reconstruction are typically implemented by machine learning models, SemCom faces model-related threats throughout its lifecycle, including training data poisoning, backdoor insertion, adversarial inputs, and privacy attacks such as model extraction and inversion. Finally, SemCom remains exposed to availability threats, including jamming and flooding, as well as computation-based denial-of-service attacks targeting the semantic processing pipeline, all of which can degrade reliability and service continuity even when the physical link remains operational.
@@ III-B1 Attacks on Probabilistic Modulation
Probabilistic modulation introduces attack surfaces that are largely unique to digital SemCom, because the transmitter does not output a single deterministic discrete representation. Instead, it learns a conditional transition distribution from source content to bits or constellation symbols using a probabilistic network and then generates the transmitted sequence by sampling from this distribution. This stochastic generation step enables distribution manipulation attacks, where an adversary perturbs the source or the semantic information to reshape the output distribution, for example by shifting probability mass toward semantically harmful alternatives, increasing ambiguity among competing hypotheses, or increasing uncertainty so that repeated sampling yields inconsistent semantic outcomes. As a result, downstream task outputs can degrade even when conventional link quality remains acceptable.
In the branch that targets discrete channel models such as the binary erasure channel and binary symmetric channel, the encoder can be viewed as a neural network-parameterized Bernoulli generator, so attacks may focus on biasing the Bernoulli probabilities of a small set of task-critical bits, thereby steering the sampled bit patterns toward attacker-preferred semantic information. In the constellation-symbol branch that resembles variational autoencoder-style modeling, a key target is the probabilistic inference network at the receiver. By biasing its inferred posterior over transmitted symbols or inducing high confidence in an incorrect hypothesis, an adversary can cause persistent semantic misinterpretation.
@@ III-B2 Attacks on Deterministic Modulation
Deterministic modulation explicitly inserts a quantizer-dequantizer interface to convert continuous semantic information into discrete representations as bits, symbols, or indices. This discretization exposes hard decision boundaries and explicit mapping rules, allowing attackers to cause large semantic distortions with small perturbations by forcing quantization or decision outcomes to change.
In scalar quantization, attackers can push selected feature elements across quantization thresholds to trigger cascaded bit changes after integer-to-bit mapping, or directly manipulate the intermediate bit stream through selective flipping, insertion, or replay before semantic decoding. In symbol quantization, remapping learned symbol-like outputs onto a predefined constellation creates constellation region vulnerabilities, where carefully designed perturbations induce systematic decision errors or configuration mismatches that consistently bias semantic interpretation. In vector quantization, semantic vectors are represented by codebook indices, so corrupting a small number of index bits can redirect decoding to different codewords and cause significant semantic deviation, while perturbations near codeword boundaries can induce nearest-codeword confusion. If endpoints are compromised, poisoning or desynchronizing the shared codebook can further lead to systematic misinterpretation. Finally, deterministic modulation typically relies on differentiable approximations to train through non-differentiable quantization, such as straight-through estimators or soft-to-hard relaxations, which can introduce a training-deployment mismatch that attackers may exploit to induce harmful discrete outcomes under hard quantization at inference.
@@ III-B3 Packet- and Protocol-Level Threats
Packet- and protocol-level threats arise primarily in digital SemCom because discrete semantic information is carried through a packet-based digital communication stack. To deliver discrete bits or symbols over practical networking procedures, digital SemCom typically segments semantic information into packets and couples them with headers, control signaling, and protocol state. This creates additional privacy leakage channels beyond the semantic message itself. Packet sizes, timing patterns, sequence structure, retransmission behavior, and explicit semantic metadata such as task identifiers, importance indicators, model or codebook version tags, and knowledge base synchronization markers can reveal sensitive intents or context through traffic analysis.
Beyond passive leakage, an adversary can actively exploit packet and protocol mechanisms to induce semantic errors. Examples include packet injection, replay, reordering, fragmentation manipulation, and selective dropping of a small subset of task-critical packets to cause disproportionate semantic distortion. Attackers can also target protocol operations such as acknowledgments, retransmissions, and rate adaptation by fabricating feedback, inducing unnecessary retransmissions or suppressing recovery, and driving the system into a suboptimal operating point. In addition, digital SemCom may depend on negotiating and synchronizing configuration state across endpoints, such as coding and modulation settings, semantic model versions, codebooks, or knowledge base updates. Manipulating these protocol exchanges can create persistent transmitter-receiver mismatch and systematic misinterpretation.
@@ IV-A Defenses Against Bit- or Symbol-Level Semantic Leakage
SEPs can be derived or predicted from channel conditions, modulation order, coding rate, and transmit power, which makes SEP-based protection implementable at the physical layer. One representative mechanism is to superpose friendly jamming, also known as artificial noise (AN), onto the transmitted signal with careful power allocation [13, 14]. The friendly jamming is designed to increase the eavesdropper’s SEP while keeping the legitimate receiver’s SEP within a tolerable range. This asymmetry can be achieved when the legitimate link has a channel advantage, such as stronger channel gain, better beamforming alignment, or more accurate channel knowledge, enabling the legitimate receiver to suppress or mitigate the interference more effectively than the eavesdropper. The legitimate receiver may further benefit from private side information that is unavailable to the eavesdropper [14]. For example, it may have access to the friendly jamming source, the jamming sequence or coding rule, and the corresponding power allocation strategy, which enables partial cancellation or mitigation of the interference before decoding. This additional capability increases the SEP gap between the legitimate receiver and the eavesdropper, improving privacy protection while preserving the intended semantic task performance.
@@
Besides physical-layer protection, encryption remains a fundamental and widely applicable tool for limiting semantic leakage in digital SemCom [15]. Since digital SemCom produces a discrete bit stream, it naturally supports conventional cryptographic primitives, such as link-layer or end-to-end authenticated encryption, to protect the semantic information against passive eavesdropping and direct feature inspection. Proper key establishment and management are essential, and encryption is often most effective when combined with integrity protection so that fabricated or manipulated semantic information is rejected rather than decoded.
@@ IV-B Defenses Against Attacks on Probabilistic Modulation
Attacks on probabilistic modulation mainly aim to reshape the learned transition distribution or to disrupt sampling behavior. A practical defense objective is therefore to make the transmitter distribution less sensitive to small perturbations and to limit unnecessary growth of predictive uncertainty. This can be supported by regularizing the probabilistic encoder to maintain stable entropy and well-calibrated outputs under channel variations, and by using perturbation-aware training so that the learned distribution remains consistent when the source or the semantic information are slightly perturbed. In addition, sampling-induced semantic inconsistency can be reduced by adopting controlled generation at deployment, such as drawing multiple samples with aggregation, applying confidence-based constraints, or switching to a more conservative selection rule when predicted uncertainty is high, thereby reducing the attacker’s ability to induce unstable task outputs through distribution manipulation.
For the bit sequence branch with Bernoulli generators parameterized by neural networks, defenses should focus on preventing the probabilities of task-critical bits from being biased by targeted perturbations. Practical measures include identifying sensitive bits or probability dimensions and applying stronger protection or redundancy to them, as well as using receiver-side consistency checks to detect abnormal probability patterns that deviate from typical operating conditions. For the constellation symbol branch, where the receiver relies on a probabilistic inference network to form posterior beliefs over transmitted symbols, robustness can be improved by hardening the inference network against adversarial channel observations, for example through adversarially augmented training and uncertainty-aware fallback or rejection, so that the network avoids producing high-confidence decisions under manipulated observations.

### v4

@@ Abstract
Semantic communication (SemCom) has emerged as a promising paradigm for future wireless networks by prioritizing task-relevant meaning over raw data delivery, thereby reducing communication overhead and improving efficiency. However, shifting from bit-accurate transmission to task-oriented delivery introduces new security and privacy risks. These include semantic leakage, semantic manipulation, knowledge base vulnerabilities, model-related attacks, and threats to authenticity and availability. Most existing secure SemCom studies focus on analog SemCom, where semantic features are mapped to continuous channel inputs. In contrast, digital SemCom transmits semantic information through discrete bits or symbols within practical transceiver pipelines, offering stronger compatibility with real-world systems while exposing a distinct and underexplored attack surface. In particular, digital SemCom typically represents semantic information over a finite alphabet through explicit digital modulation, following two main routes: probabilistic modulation and deterministic modulation. These discrete mechanisms and practical transmission procedures introduce additional vulnerabilities affecting bit- or symbol-level semantic information, the modulation stage, and packet-based delivery and protocol operations. Motivated by these challenges and the lack of a systematic analysis of secure digital SemCom, this paper provides a structured review of the area. Specifically, we review SemCom fundamentals and clarify the architectural differences between analog and digital SemCom. We then summarize threats shared by both paradigms and organize the threat landscape specific to digital SemCom, followed by a discussion of potential defenses. Finally, we outline open research directions toward secure and deployable digital SemCom systems.
@@ I Introduction
However, the characteristics that make SemCom efficient and task-oriented also raise new concerns for security and privacy. Since SemCom often transmits task-oriented representations, the transmitted content may inadvertently reveal sensitive information. In particular, an eavesdropper may not need to reconstruct the source data to obtain useful information, because semantic information can already enable inference of private attributes. Beyond passive leakage, SemCom is also vulnerable to semantic manipulation. In such attacks, an adversary perturbs the source or semantic information to distort the recovered meaning and mislead downstream tasks, potentially causing misinformation, unsafe decisions, or service disruption. In addition, SemCom commonly relies on data-driven models and knowledge bases, expanding the attack surface beyond the physical channel. Adversaries may compromise the learning pipeline through data poisoning or backdoor insertion, exploit model-related vulnerabilities such as model extraction and inversion, or manipulate the knowledge base to bias semantic interpretation and induce systematic errors. Addressing these security and privacy challenges is therefore essential for practical deployment.
@@
In practice, analog and digital SemCom differ in how semantic information is transmitted over the wireless interface, leading to distinct attack surfaces and design constraints. Analog SemCom typically maps semantic features directly to continuous channel inputs and relies on joint source-channel behavior. Digital SemCom, in contrast, transmits semantic information through discrete bits or symbols within practical digital transceiver pipelines. In such designs, semantic information is mapped into a finite alphabet via explicit digital modulation, following two main routes: probabilistic modulation, which learns a conditional distribution and generates discrete symbols using neural network models, and deterministic modulation, which integrates explicit quantization. Meanwhile, this discrete interface also facilitates the integration of cryptographic primitives. This pipeline introduces additional vulnerabilities, including bit- or symbol-level semantic leakage and manipulation. Moreover, digital SemCom also faces threats such as vulnerabilities within probabilistic or deterministic modulation components and attacks arising from packet-based delivery and protocol operations. These characteristics, together with the underexplored security risks in digital SemCom, motivate our dedicated focus on secure digital SemCom to provide a structured understanding of its threat landscape and corresponding defenses.
@@ II Background Knowledge of Semantic Communication
Figure 2: The detailed architecture of a digital SemCom system with explicit probabilistic or deterministic modulation and packet-based delivery. The JSC encoder/decoder denotes the joint source and channel coding/decoding module. Packet channel is a logical abstraction over the physical channel, capturing packetization and protocol effects.
@@
The transmitter then modulates the semantic information into a transmitted signal, which propagates through the wireless channel and experiences fading, noise, and interference. At the receiver, the received signal is demodulated, and channel decoding mitigates channel distortions to recover reliable semantic information. The semantic decoder then reconstructs the intended meaning or directly outputs the task result for the information destination. The information destination is defined at the receiver and reflects the target task. Throughout this process, a knowledge base can be maintained to store domain knowledge and contextual information, helping the transmitter and receiver align semantic interpretation and improve semantic consistency under limited communication resources.
@@
In probabilistic modulation [11], the modulator is modeled as a stochastic mapping implemented by a neural encoder that outputs a conditional distribution over a finite bit or symbol alphabet. The transmitted sequence is generated by sampling from this learned distribution, while the demodulator is realized by a probabilistic inference network that estimates the posterior distribution of the transmitted bits or symbols given the received signal. A key challenge is that end-to-end training must handle discrete random variables, making posterior inference and gradient estimation nontrivial. As a result, probabilistic digital SemCom often adopts techniques such as variational inference, score-function estimators, reparameterization-based approximations, and continuous relaxations such as Gumbel-Softmax. It also accounts for potential mismatch between relaxed training and discrete inference at deployment.
@@ III-A Common Challenges in Analog and Digital Semantic Communication
Semantic leakage and inference attacks are among the most fundamental privacy challenges in SemCom. Because the transmitted message is task-oriented semantic information, intercepted semantic information can already reveal sensitive attributes, intents, or environmental states. The risk is amplified when the adversary combines the intercepted content with side information or public background knowledge. SemCom is also vulnerable to semantic manipulation, where an attacker introduces carefully designed perturbations to distort the recovered meaning and mislead downstream tasks. Such attacks may not be reflected by conventional link-level metrics while still causing severe semantic errors. They also raise concerns about semantic integrity and source authenticity, since the receiver must determine whether the recovered meaning is reliable and originates from a legitimate transmitter rather than from fabricated or manipulated semantic information.
Another set of shared challenges arises from the supporting context and learning components. Many SemCom designs rely on shared context or knowledge base to align interpretation across endpoints. This introduces risks such as knowledge poisoning, manipulation, and inconsistencies between transmitter and receiver knowledge states that can systematically bias semantic interpretation. Moreover, since semantic extraction and reconstruction are typically implemented by machine learning models, SemCom is exposed to model-related threats throughout its lifecycle. These threats include training data poisoning, backdoor insertion, and privacy attacks such as model extraction and inversion. Finally, SemCom remains exposed to availability threats, including jamming and flooding, as well as computation-based denial-of-service attacks targeting the semantic processing pipeline. These attacks can degrade reliability and service continuity even when the physical link remains operational.
@@ III-B1 Attacks on Probabilistic Modulation
Attacks on probabilistic modulation are more pronounced in digital SemCom because the transmitted representation is generated by sampling from a learned conditional distribution. An adversary can target the distribution parameters to shift probability mass or alter distribution concentration, which biases the sampling behavior toward realizations that cause errors in the semantic decoding outcome, possibly in a targeted manner.
In the branch that targets discrete channel models such as the binary erasure channel (BEC) and binary symmetric channel (BSC), the encoder acts as a neural-network-parameterized Bernoulli generator whose outputs determine per-bit sampling probabilities. A primary threat is probability biasing, where an adversary induces small but structured perturbations to the source or intermediate semantic information, and thereby causes shifts in a subset of Bernoulli parameters. As a result, the sampled bit patterns become more likely to trigger semantic decoding errors at the receiver.
In the constellation-symbol branch based on variational autoencoder (VAE)-style modeling, a key target is the probabilistic inference network at the receiver. By manipulating the inferred posterior so that probability mass concentrates on an incorrect hypothesis, or by inducing overconfident posteriors around the wrong mode, an adversary can cause systematic misinterpretation of semantic information.
@@ III-B2 Attacks on Deterministic Modulation
Deterministic modulation inserts a quantizer-dequantizer interface to convert continuous semantic information into discrete representations such as bits, symbols, or indices. This discretization exposes hard decision boundaries and fixed mapping rules. As a result, small but structured perturbations near quantization or decision boundaries can be amplified into discrete changes that propagate through decoding and lead to erroneous semantic decoding outcomes.
In scalar quantization, an adversary can perturb selected feature elements to push them across quantization thresholds. Depending on the integer-to-bit mapping, such threshold crossings can induce multiple bit-level changes and thereby increase the likelihood of semantic decoding errors. In symbol quantization, remapping learned symbol-like outputs onto a predefined constellation introduces decision regions that can be exploited by carefully designed perturbations. Such perturbations can shift received points across region boundaries and cause systematic symbol decision errors.. In vector quantization with a shared codebook, semantic vectors are represented by codebook indices before bit mapping and conventional digital modulation. Perturbations near codeword boundaries can induce nearest-codeword confusion, and corruption of index bits can redirect decoding toward unintended codewords. Because vector quantization relies on a shared codebook, codebook integrity and consistency between endpoints are critical. Compromising the codebook through poisoning, unauthorized updates, or desynchronization can lead to systematic semantic misinterpretation.
@@ III-B3 Packet- and Protocol-Level Threats
Packet- and protocol-level threats are often more salient in digital SemCom because packet-level transport inherently operates on discrete representations. In addition, digital SemCom may introduce semantic-related signaling and coordination state at the packet or control level. This introduces privacy leakage beyond the semantic payload through observable traffic features such as packet sizes, timing, sequencing, and retransmission patterns. Additional leakage can arise from explicit semantic-related metadata carried in packets or control messages, including task identifiers, importance indicators, and version tags for models, codebooks, or knowledge bases. This metadata may enable an adversary to infer sensitive intent or context via traffic analysis.
Beyond passive leakage, adversaries can exploit packet and protocol mechanisms to induce semantic errors. General primitives such as injection, replay, reordering, fragmentation manipulation, and selective dropping can be amplified when semantic-aware packetization makes a small subset of high-impact packets dominate the semantic decoding outcome. Manipulating acknowledgments, retransmissions, and rate adaptation can further degrade semantic decoding quality, especially when feedback or control signaling is unauthenticated or weakly protected.
@@ IV-A Defenses Against Bit- or Symbol-Level Semantic Leakage
SEPs can be derived or predicted from channel conditions, modulation order, coding rate, and transmit power, which makes SEP-based protection implementable at the physical layer. One representative mechanism is to superpose friendly jamming, also known as artificial noise (AN), onto the transmitted signal with careful power allocation [13, 14]. The friendly jamming is designed to increase the eavesdropper’s SEP while keeping the legitimate receiver’s SEP within a tolerable range. This asymmetry can be achieved when the legitimate link has a channel advantage, such as stronger channel gain, better beamforming alignment, or more accurate channel knowledge, enabling the legitimate receiver to suppress or mitigate the interference more effectively than the eavesdropper. The legitimate receiver may further benefit from private knowledge that is unavailable to the eavesdropper [14]. For example, it may have access to the friendly jamming source, the jamming sequence or coding rule, and the corresponding power allocation strategy, which enables partial cancellation or mitigation of the interference before decoding. This additional capability increases the SEP gap between the legitimate receiver and the eavesdropper, improving privacy protection while preserving the intended semantic task performance.
@@
Besides physical-layer protection, encryption remains a fundamental and widely applicable tool for limiting semantic leakage in digital SemCom [15]. Since digital SemCom produces a discrete bit stream, conventional cryptographic primitives can be readily integrated. For example, link-layer or end-to-end authenticated encryption can protect semantic information against passive eavesdropping and direct feature inspection. Proper key establishment and management are essential. Encryption is often most effective when combined with integrity protection, so that fabricated or manipulated semantic information is rejected rather than decoded.
@@ IV-B Defenses Against Attacks on Probabilistic Modulation
Attacks on probabilistic modulation mainly aim to reshape the learned conditional distribution or to disrupt sampling behavior. A practical defense objective is therefore to make the transmitter distribution less sensitive to small perturbations in the source or intermediate semantic information. It should also prevent attack-induced inflation of predictive uncertainty or overconfident concentration that can amplify semantic errors. This can be supported in two ways. First, by regularizing the probabilistic encoder so that its output distribution remains reasonably well calibrated and exhibits controlled uncertainty under channel variations. Second, by using perturbation-aware training so that the learned distribution changes smoothly when the source or the semantic information is slightly perturbed. As a result, the distribution avoids large probability-mass shifts and abrupt concentration changes. In addition, sampling-induced inconsistency in the semantic decoding outcome can be reduced by adopting controlled sampling at deployment. For example, the receiver can draw a limited number of samples and aggregate them, enforce confidence- or entropy-based constraints, or switch to more conservative selection rules when predicted uncertainty is high. These measures reduce the adversary’s ability to induce unstable task outputs through distribution manipulation while introducing extra latency only when predicted risk is high.
For the bit-sequence branch with Bernoulli generators parameterized by neural networks, defenses should focus on limiting how much targeted perturbations can bias a subset of Bernoulli parameters and how strongly such biases propagate to erroneous decoder output and semantic decoding errors. Practical measures include identifying Bernoulli probability dimensions that are unusually sensitive or semantically critical. Stronger protection or redundancy can then be applied to these dimensions to reduce the effect of probability biasing on sampled bit patterns. For the constellation-symbol branch, the receiver typically relies on an explicit probabilistic inference network to form posterior beliefs over transmitted symbols. Robustness can be improved by hardening the posterior inference network against adversarial channel observations through adversarially augmented training and by introducing uncertainty-aware fallback or rejection mechanisms. This helps reduce the likelihood that the receiver concentrates posterior mass on incorrect hypotheses or produces overconfident posteriors around a wrong mode under manipulated observations.